OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the third equation. 2x – 3y + z = –4 2x – 2y + z = –1 x – 2y + 3z = –6

OpenStudy (anonymous):

Choices:

OpenStudy (anonymous):

5x + 7y = 6 y – 4 = –1 –y – 1 = –4 –5x + 4y = –3 y – 5z = 8 2y – 5z = 11 –7y + 7z = 8 –6y + 7z = 11

OpenStudy (anonymous):

If x was isolated in the third equation, you would get x=2y-3z-6, use that to 'substitute' back into the first two equation to get your answer. \[2*(2y-3z-6) -3y +z = -4\] \[2*(2y-3z-6) - 2y +z = -1\]

OpenStudy (anonymous):

4y - 6z - 12 - 3y + z = -4

OpenStudy (anonymous):

y - 5z = 8

OpenStudy (anonymous):

4y - 6z - 12 - 2y + z = -1

OpenStudy (anonymous):

2y - 5z = 11

OpenStudy (anonymous):

so it's c?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!